OpenStudy (texaschic101):

you see my avatar...this is what my real face looks like when I am viewing this problem...except I am not yellow

OpenStudy (texaschic101):

My grandson is as many days as my son is in weeks. My grandson is as many months as I am in years. We are all three, when added together, 120 years. How old am I , in years ?

OpenStudy (texaschic101):

m + g + s = 120 thats all I got

OpenStudy (texaschic101):

I have no idea

OpenStudy (texaschic101):

why ?

OpenStudy (texaschic101):

I have one equation.....I don't know where to go from there

OpenStudy (mathstudent55):

Let g = age of granson in days Let s = age of son in weeks Let m = my age in years Let's convert all ages to days. m + g + s = 120(265.25)

OpenStudy (texaschic101):

wait...where did you get the 265.25 ?

OpenStudy (mathstudent55):

There are 365.25 days in a year.

OpenStudy (texaschic101):

oh....but you got 265.25...why are you subtracting 100

OpenStudy (mathstudent55):

typo

OpenStudy (texaschic101):

ok...I got you

OpenStudy (mathstudent55):

The problem is you need to make an assumption of how many days per month.

OpenStudy (texaschic101):

30 on average

OpenStudy (mathstudent55):

Let's do it your way with everything in years. Your first equation is good.

OpenStudy (mathstudent55):

m + g + s = 120 Keep in mind that m, s, and g are all in years.

OpenStudy (texaschic101):

ok

OpenStudy (mathstudent55):

"My grandson is as many days as my son is in weeks." We need to change the above statement into an equation.

OpenStudy (mathstudent55):

How many days are there in g years? How many weeks are there in s years?

OpenStudy (texaschic101):

365.25 days 52 weeks/year

OpenStudy (texaschic101):

365.25g = 52s ??

OpenStudy (mathstudent55):

Ok, so g years = 365.25g days s years = 52s weeks 365.25g = 52s You are correct.

OpenStudy (mathstudent55):

Now we have tw equations. We need one more.

OpenStudy (texaschic101):

do I solve for g ?

OpenStudy (mathstudent55):

Not yet. Let's get our 3rd equation so we end up with a system of three equations in three unknowns. "My grandson is as many months as I am in years."

OpenStudy (texaschic101):

12g = m ??

OpenStudy (mathstudent55):

g years = 12g months m years = m years 12g = m Great. You got it.

OpenStudy (mathstudent55):

m + g + s = 120 365.25g = 52s 12g = m

OpenStudy (texaschic101):

ok... m + g + s = 120 365.25g = 52s 12g = m so I sub 12g in for m in the 1st equation ?

OpenStudy (mathstudent55):

Yes

OpenStudy (texaschic101):

12g + g + s = 120 13g + s = 120 13g = 120/s

OpenStudy (mathstudent55):

and then the first 2 eqs are a system of 2 eqs in 2 unknowns.

OpenStudy (mathstudent55):

Careful: 13g + s = 120 13g = 120 - s

OpenStudy (mathstudent55):

Now solve for g: 13g = 120 - s

OpenStudy (texaschic101):

13g = 120 - s g = (120 - s)/13 g = 9.23 - s/13

OpenStudy (mathstudent55):

Now use the second equation: 365.25g = 52s Substitute g with 9.23077 - s/13

OpenStudy (texaschic101):

365.25 (9.23077 - (s/13) = 52s

OpenStudy (mathstudent55):

365.25(9.23077 - s/13) = 52s

OpenStudy (mathstudent55):

Good. Multiply out the left side.

OpenStudy (mathstudent55):

3371.538 - 28.096s = 52s

OpenStudy (texaschic101):

3371.5387425 - 365.25s/13 = 52s do I multiply the equation by 13 ?

OpenStudy (texaschic101):

oops

OpenStudy (mathstudent55):

Don't erase. You're ok.

OpenStudy (texaschic101):

3371. 538 = 52s + 28.096s am I on the right track

OpenStudy (mathstudent55):

great

OpenStudy (mathstudent55):

Add the right side. Divide both sides by the number multiplying s.

OpenStudy (texaschic101):

3371.538 = 80.096s 3371.538/80.096 = s 42.0937 = s

OpenStudy (mathstudent55):

Great. You got one of them.

OpenStudy (mathstudent55):

Now use s here to get g. 365.25g = 52s

OpenStudy (texaschic101):

13g = 120 - s 13g = 120 - 42.0937 13g = 77.9603 g = 77.9603/13 g = 5.9928

OpenStudy (texaschic101):

oops...I did the wrong one

OpenStudy (mathstudent55):

Excellent. Finally, use 12g = m to find m.

OpenStudy (mathstudent55):

It makes no difference which eq you use as long as it's a correct eq.

OpenStudy (texaschic101):

12(5.9928) = m 71.9136 = m

OpenStudy (texaschic101):

do I round any numbers.....would m be 72

OpenStudy (mathstudent55):

Ok, so in years the ages are: Grandparent (me) m = 71.9 Son s = 42.1 Grandson g = 5.99 Right?

OpenStudy (texaschic101):

m + g + s = 120 71.9 + 5.99 + 42.1 = 120 119.99 = 120 rounded 120 = 120 yes.... wow, I am 72...I feel that old

OpenStudy (texaschic101):

thank you soooooooo much

OpenStudy (mathstudent55):

Let's see if our answers make sense with the given info. "My grandson is as many days as my son is in weeks." 5.99(365.25) = 2188 days old for grandson 42.1(52) = 2189 weeks old for son This checks out.

OpenStudy (texaschic101):

your the best

OpenStudy (mathstudent55):

"My grandson is as many months as I am in years." 5.99(12) = 72 months for grandson 72 grandparent's age in years This also checks out.

OpenStudy (texaschic101):

you should be a teacher

OpenStudy (mathstudent55):

Finally, he one you already checked: "We are all three, when added together, 120 years." 71.9 + 5.99 + 42.1 = 120 This checks out, too.

OpenStudy (mathstudent55):

You did a great job. Thanks for you nice comment.

OpenStudy (texaschic101):

you did a better job.....thanks again

OpenStudy (mathstudent55):

You can now fix your avatar to a nice smile! You're welcome.

OpenStudy (texaschic101):

lol...I might just do that